OpenStudy (anonymous):

sqrt 45x/16y^4 divide and simplify

jhonyy9 (jhonyy9):

3sqrt5x ------- 4y^2

OpenStudy (anonymous):

\[\frac{\sqrt{45x}}{16y^4}=\frac{\sqrt{9*5x}}{16y^4}=\frac{3\sqrt{5x}}{16y^4}\]